Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.50 coming 3 years from today. The dividend should grow rapidly-at a rate of 40% per year-during Years 4 and 5; but after Year 5, growth should be a constant 5% per year. If the required return on Computech is 14%, what is the value of the stock today? Round your answer to the nearest cent. Do not round your intermediate calculations.
Answer #1

D3=1.5

D4=(1.5*1.4)=2.1

D5=(2.1*1.4)=2.94

Value after year 5=(D5*Growth rate)/(Required rate-Growth Rate)

=(2.94*1.05)/(0.14-0.05)

=34.3

Hence value of stock today=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.5/1.14^3+2.1/1.14^4+2.94/1.14^5+34.3/1.14^5

=$21.60(Approx).